# Arithmetic in Transformers Explained

### v9

@@ Abstract
While recent work has shown transformers can learn addition, previous models exhibit poor prediction accuracy and are limited to small numbers. Furthermore, the relationship between single-task and multitask arithmetic capabilities remains unexplored. In this work, we analyze 44 autoregressive transformer models trained on addition, subtraction, or both. These include 16 addition-only models, 2 subtraction-only models, 8 “mixed” models trained to perform addition and subtraction, and 14 mixed models initialized with parameters from an addition-only model. The models span 5- to 15-digit questions, 2 to 4 attention heads, and 2 to 3 layers. We show that the addition models converge on a common logical algorithm, with most models achieving >99.999% prediction accuracy. We provide a detailed mechanistic explanation of how this algorithm is implemented within the network architecture. Subtraction-only models have lower accuracy. With the initialized mixed models, through parameter transfer experiments, we explore how multitask learning dynamics evolve, revealing that some features originally specialized for addition become polysemantic, serving both operations, and boosting subtraction accuracy. We explain the mixed algorithm mechanically. Finally, we introduce a reusable library of mechanistic interpretability tools to define, locate, and visualize these algorithmic circuits across multiple models.
Machine Learning, ICML
11footnotetext: Corresponding author: philipquirkenz@gmail.com
@@ 1 Introduction
Figure 1: An overview of our methodology: (1) We trained an accurate 5-digit integer addition model. (2) We reverse-engineered the model to find the algorithms that were implemented to perform addition. (3) We inserted the addition model into a new model, by copying the weights of the attention heads and MLPs (in brown) into the larger model during initialization. (4) We then train the new model on 80% subtraction and 20% addition questions. (5) We find that the resulting model predicts accurately and reuses the inserted addition circuits for both addition and subtraction questions.
Transformer-based large language models (LLMs) are powerful (Barak et al., 2022) yet largely inscrutable due to their complex, nonlinear interactions in dense layers within high-dimensional spaces. Given this complexity, their deployment in critical settings (Zhang et al., 2022) highlights the need to understand their behavior. Hendrycks & Mazeika (2022) argue that making these models interpretable is key to their safe use.
Previous work by (Quirke & Barez, 2024) demonstrated that transformers can learn addition on small numbers, but their models had poor accuracy and were limited to a single operation. Understanding how transformers can overcome these limitations to perform more accurate mathematical computations across multiple operations is crucial for several reasons. First, it provides insights into how neural networks can learn and combine multiple precise algorithms, a key capability for more complex reasoning tasks. Second, studying how models transition from single-task to multitask competence while maintaining accuracy reveals important principles about knowledge transfer and representation sharing in neural networks.
Understanding how models arrive at their predictions is critical for diagnosing errors, enhancing safety, and gaining fundamental insights into modern neural networks. Mechanistic interpretability provides a pathway toward this understanding by identifying the “circuits” responsible for particular computations within a network’s layers and attention heads.
In this work, we use multidigit arithmetic as a revealing testbed for mechanistic interpretability. We train and analyze 44 autoregressive transformer models on addition, subtraction, or both, covering question lengths of 5 to 15 digits and varying in their number of layers (2–3) and attention heads (2–4). Our experiments encompass three categories of models: addition-only, subtraction-only, and “mixed” models trained to handle both operations. Through a systematic investigation of these networks, we show that addition-only models converge on a highly accurate, shared algorithm, yielding near-perfect performance. We mechanistically explain the algorithm. Subtraction-only models, in contrast, exhibit lower accuracy.
To explore how multitask learning reshapes these internal algorithms, we then conduct parameter transfer experiments using three distinct approaches, each involving the initialization of mixed models with parameters from addition-only networks. Interestingly, one of these approaches boosts the subtraction accuracy beyond what is achieved by mixed models trained from scratch, indicating that features specialized for addition can be effectively repurposed for subtraction. We provide a detailed mechanistic explanation of the mixed algorithm, illustrating how certain addition circuits become “polysemantic” under multitask learning, driving behavior for both operations.
A further contribution of our study is the development of a reusable library of mechanistic interpretability tools. This library enables researchers to define and search for algorithmic features, visualize network structures and feature interactions, and conduct cross-model comparisons efficiently.
Overall, our results demonstrate the power of mechanistic interpretability in shedding light on how transformers learn and reuse algorithmic circuits across related tasks, paving the way for deeper understanding and more reliable deployment of large language models in diverse domains.
@@ 2 Related Work
Mechanistic interpretability aims to reverse engineer neural networks to find interpretable algorithms that are implemented in a model’s weights (Olah et al., 2020). Mathematical frameworks (Elhage et al., 2021a) explain how transformer attention heads can work with each other to implement complex algorithms.
Causal Scrubbing (Jenner et al., 2023) recommends explaining a model algorithm by documenting a low-level computation graph, mapping from the graph to the model nodes that implement the computation, and performing experimentation verification. Investigative techniques such as ablation interventions, activation unembeddings (nostalgebraist, 2020), and sparse autoencoders (Nanda, 2023; Cunningham et al., 2023), underpinned by the more theoretical frameworks (Elhage et al., 2021b; Geva et al., 2022), provide tools to help confirm a mapping.
Investigating pre-trained LMs on Arithmetic. Even though basic arithmetic can be solved following a few simple rules, pre-trained LMs often struggle to solve simple math questions (Hendrycks et al., 2021). Causal mediation analysis (Stolfo et al., 2023) has been used to investigate how large pre-trained LMs like Pythia and GPT-J performed addition to solve word problems. It is also possible to improve a model’s arithmetic abilities used supervised fine tuning - including enriched training data (Liu & Low, 2023).
Studying Toy Models for Arithmetic. Applying mechanistic interpretability to toy transformers can help to better isolate clear, distinct circuits given the highly specific experimental setup for the model studied (Nanda et al., 2023). Quirke & Barez (2024) detailed a 1-layer, 3-head transformer model that performs 5-digit addition, showing it failed on edge cases (e.g. “77778+22222=100000” where a “carry 1” cascades through 4 digits), incorrectly predicting 1% of questions.
Many natural prediction problems decompose into a finite set of knowledge and skills that are “quantized” into discrete chunks (quanta) (Michaud et al., 2023). Models must learn these quanta to reduce loss. Understanding a network reduces to enumerating its quanta. Other studies (Schaeffer et al., 2023) prove useful ways to measure quanta in mathematical prediction problems.
@@ 3 Methodology
Transformer models learn addition algorithms different from traditional human methods. We define an alternative, mathematically-equivalent framework for addition and demonstrate our models implements this approach.
@@ 3.1 Mathematical Framework
Consider the task of adding two $n$-digit numbers together. We define the first number (or addend) as $D=\{D_{n-1},D_{n-2},\ldots,D_{0}\}$ and the second number as $D^{\prime}=\{D^{\prime}_{n-1},D^{\prime}_{n-2},\ldots,D^{\prime}_{0}\}$ and the answer (or sum) as $A=\{A_{n},A_{n-1},\ldots,A_{0}\}$. Figure 2 shows an illustrative example.
@@
Figure 2: For 5-digit addition, our model has 12 input (question) and 6 output (answer) token positions. We name the question tokens D4, …, D0, and D’4, …, D’0 and the answer tokens A5, …, A0. For n-digits, we use the terms $D_{n}$, $D^{\prime}_{n}$and $A_{n}$.
We adopt the framework from Quirke & Barez (2024) for our model’s addition process, where they break down the task of multi-digit addition into an addition subtask and a carry subtask. Firstly, the “Simple Addition” subtask $A_{n}.SA$ naively calculates the sum of corresponding digit pairs in the addends:
$A_{n}.SA=(D_{n}+D^{\prime}_{n})\mod 10$ (1)
The “Simple Carry” subtask $A_{n}.SC$ then determines whether the addition between the two digits creates a carry bit:
$A_{n}.SC=\begin{cases}1&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ 0&\text{otherwise}.\end{cases}$ (2)
By combining these two subtasks, the model and framework from Quirke & Barez (2024) is capable of handling simple addition and carry bits generated directly from digit pair addition. However, it struggles with problems with ’cascading carry’ bits, where a carry bit from one digit position propagates to the next.
Consider “00144+00056=000210”. Adding 4+5 in the tens position doesn’t generate a carry bit directly, but a carry bit propagates from the ones position. A model only summing digit pairs and their direct carry bits would fail, producing an incorrect result like “00144+50006=000110”. The Quirke & Barez (2024) 1-layer model could cascade carry bits across two digits, but not three or more.
3.2 Extending the Mathematical Framework
@@
To solve $n$-digit addition with high accuracy, we need to extend the basic framework to handle carry bits that cascade through multiple digits. Consider the problem ”44444+55556=100000”. For an autoregressive model to correctly predict the first answer digit as ”1”, it needs to track how the carry bit cascades from the rightmost digit all the way to the leftmost position in a single forward pass. This is particularly challenging for an autoregressive model that predicts answer tokens from left to right.
To solve this, we extend the mathematical framework with two key subtasks.
First, the digit-level subtask ”Sum Type” $A_{n}.ST$ analyzes each pair of digits and classifies their sum as either definitely causing a carry, possibly causing a carry, or definitely not causing a carry. This is implemented by a TriCase operation that maps each digit pair ($D_{n}$ and $D^{\prime}_{n}$), TriCase into one of three possible states:
$\begin{split}A_{n}.ST&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}ST10&\text{if }(D_{n}+D^{\prime}_{n})\geq 10\text{ (definite carry)}\\ ST9&\text{if }(D_{n}+D^{\prime}_{n})=9\text{ (possible carry)}\\ ST8&\text{if }(D_{n}+D^{\prime}_{n})\leq 8\text{ (no carry)}\\ \end{cases}\end{split}$ (3)
The key insight is that while some digit pairs definitely generate a carry (like 5+6), and others definitely don’t (like 2+3), pairs that sum to 9 (like 4+5) may or may not generate a carry depending on whether there’s an incoming carry bit from the next lower digit.
Then, we introduce the ”Sum Validation” subtask $A_{n}.SV$ that processes these carry states sequentially to determine whether any ”possible carries” become ”definite carries”. This subtask is implemented using a TriAdd operation that examines each digit’s sum type along with the sum type of the next lower digit:
$\begin{split}A_{n}.SV&=\underbrace{\text{TriAdd}}_{(A_{n}.ST,A_{n-1}.ST)}=\\ &\begin{cases}ST10&\text{if }A_{n}.ST=ST10\text{ or }\\ &\quad(A_{n}.ST=ST9\text{ and }\\ &\quad A_{n-1}.ST=ST10),\\ ST8&\text{otherwise.}\end{cases}\end{split}$ (4)
Revisiting the problem ”44444+55556=100000”, the model can accurately determine whether the first answer digit A5 should be 1 or 0 by examining all digit pairs in sequence. This chain of operations uses TriAdd to process the Sum Types from least to most significant digits:
$A1.SV$ $=\text{TriAdd}(A1.ST,A0.ST)$
$A2.SV$ $=\text{TriAdd}(A2.ST,A1.SV)$
$A3.SV$ $=\text{TriAdd}(A3.ST,A2.SV)$
$A4.SV$ $=\text{TriAdd}(A4.ST,A3.SV)$
$A5$ $=(A4.SV=ST10)$
Most importantly, by using this cascading calculation to calculating A5, the model also simulatenously calculates the accurate carry bits for each answer digit. The carry bits are simply the SV values we computed along the way—we already have A3.SV for A4’s carry bit, A2.SV for A3’s carry bit, and so on.
Within our extended framework, the model only requires the subtasks $A_{n}$.SA, $A_{n}.ST$ and $A_{n}.SV$ to accurately perform $n$-digit addition accurately. Some models may also use the redundant $A_{n}$.SC subtask.
Figure 3 shows how our model’s algorithm uses SA, SC, $ST$ and $SV$ to perform addition. This algorithm is like the 99% accurate Quirke & Barez (2024) algorithm, but contains an additional circuit (the shaded boxes) to calculate “cascading carry one” data to solve $n$-digit addition with >99.999% accuracy.
Figure 3: To predict with >99.999% accuracy, the addition algorithm first calculates “carry one” values $(A_{n}.ST)$, combining them into “cascading carry one” values $(A_{n}.SV).$ At the “+” token, $A_{n-1}.SV$ gives the first answer digit as 1 or 0. The other answer digits are calculated by combining “base add” $(A_{n}.BA)$ and “carry one” $(A_{n}.SC)$ calculations with the pre-calculated $A_{n}.SV$ values.
3.3 Techniques
To undetsand how the model implements arithmetic operations, we analyzed both attention heads and MLP layers at each token position. We define a node as either an attention head or MLP layer at a specific token position (e.g., the first attention head in the first layer processing the third digit). We employed four techniques to understand each node’s function:
  1. 1.
    Intervention Ablation. To determine the role of a node’s output, we replace its output with the mean vector of its outputs across a batch, then measured how this impacted downstream predictions. For each subtask, we developed automated intervention ablation tests across $n$-digits to verify whether a node implemented that subtask.
  2. 2.
    Attention Patterns. For each attention head node, we examined which input tokens received significant attention (defined as post-softmax attention score $>0.01$). This revealed which parts of the input each attention head considered relevant.
  3. 3.
    Principal Component Analysis (PCA). We applied PCA to node outputs, especially when investigating attention heads that our framework suggests might have descrete states (e.g. tri-state or bi-state outputs).
  4. 4.
    Question Complexity. We categorized questions by computational complexity (App. G). For addition, categories S0-S4 reflect how many sequential digits a “carry one” operation cascades through. For subtraction, categories M0-M4 reflect a similar cascading of ”borrow one” operations. We analyzed which nodes were necessary for correct predictions in each category.
@@ 4 Experiments
4.1 Training a Five-Digit Addition Model
Previous work by Quirke & Barez (2024) achieved $\sim$ 99% accuracy on 5-digit addition using a 1-layer addition model. To improve upon this, we systematically explored different model architectures and found that a transformer with 2 layers and 3 attention heads per layer was the minimal configuration needed to achieve >99.999% accuracy (see App. D for our exploration of alternative architectures, and App. B for model configuration details.). This 2-layer architecture doubles the model’s computational capacity and enables cross-layer attention composition for more sophiscated algorithms (Elhage et al., 2021b).
We trained our model to perform 5-digit addition using an architecture with 2 layers and 3 attention heads per layer. The model’s vocabulary consisted of 14 tokens: the digits 0-9 plus the mathematical operators +, -, =, *, and /. We used a batch size of 64, learning rate of 0.00008, and weight decay of 0.1.
To ensure robust performance on edge cases, we trained on an infinite dataset that was deliberately enriched with rare numerical combinations. The loss function was defined as the mean negative log likelihood across all output tokens. After 30,000 training batches, the model achieved a final training loss of approximately $2.3\times 10^{-8}$, and subsequent testing confirmed accuracy exceeding 99.999%. For comprehensive training details and results, see App. C and Tab. 6.
Table 1: For a sample model, all nodes used in predictions are shown by token position (horizontally) and model layer (vertically), detailing the answer digits they impact. Here, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5. For all addition and mixed models studied, before the ”=” token, each node often calculates data used to predict multiple answer digits. After the ”=” token, all nodes in a given token position are used to predict a single answer digit.
(P6) (P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17)
D’4 D’1 D’0 = + A5 A4 A3 A2 A1
L0H0 A4 A2 A1
L0H1 A5 A5..3 A5 A3 A0
L0H2 A5 A5..1
L0MLP A5..2 A4 A3 A2 A1 A0
L1H0
L1H2 A4 A3 A2
L1MLP A1 A0
4.2 Investigating Five-Digit Addition
Through ablation experiments, we found that the model relies on just 36 nodes distributed across nine token positions (Figs 7 and 8). We analyzed how these nodes affect computation complexity and answer accuracy through ablation studies (see Figs 6 and Table 1).
Core Algorithm Discovery. The model’s most challenging task is predicting A5, the first answer digit, at position P11. This digit, which can only be 0 or 1, depends on potential carry-bit cascades through multiple positions (e.g., in 55555+44445=100000). After two unsuccessful algorithm hypotheses (detailed in App.H andI), we discovered the model’s fundamental mechanism: it uses a minimal representation of carry information through what we call the TriCase quanta. The model computes $A_{n}.ST$ using bigram operations that map digit pairs to result tokens (e.g., ”6” + ”7” = ST10) in positions P8 to P11.
For a specific 5-digit addition model instance, we mapped the algorithm to individual nodes and verified each node’s role using ablation intervention. Figure 3 diagrams the algorithm, with node details in App. J. The model adheres to all known constraints and achieves >99.999% accuracy using 36 nodes, as compared to the 1-layer model uses 21 nodes to achieve 99% accuracy.
Table 2: All addition models studied implement our addition algorithm. The algorithm SA, SC and $ST$ subtasks all exist for each digit and in appropriate token positions. For a sample model, this map shows the subtask locations. Interestingly, here each SA subtask is shared across two attention heads.

### v10

# Understanding Addition and Subtraction in Transformers

@@ Abstract
Transformers are widely deployed in large language models (LLMs), yet most models still fail on basic arithmetic tasks such as multidigit addition. In contrast, we show that small transformers trained from scratch can solve n-digit addition and subtraction with 99.999% accuracy. Building directly on prior work that uncovered addition circuits, we extend the analysis to subtraction and present a unified mechanistic account based on cascading carry and borrow circuits. Using a suite of 49 trained models, we apply systematic ablations and node-level constraints to validate the learned mechanisms, and release a reproducible interpretability toolkit for studying arithmetic circuits. Finally, surveying 180 publicly available LLMs, we find that only 7% can reliably perform addition, underscoring the gap between specialized small models and general-purpose LLMs. Our results show that arithmetic can be implemented exactly by tiny transformers, offering a tractable case study for mechanistic interpretability and a cautionary contrast with the persistent arithmetic failures of much larger models.
@@ 1 Introduction
Large language models (LLMs) achieve remarkable performance across diverse tasks, but still struggle with basic arithmetic. In a survey of 180 LLMs ranging from 1B to 405B parameters, 7% can reliably perform addition. The few that succeed typically rely on external tools, suggesting that the underlying models have not truly learned arithmetic (See App. E). In contrast, we show that transformers trained from scratch can master both addition and subtraction with near-perfect accuracy - without using external tools.
Figure 1: Our n-digit addition algorithm is mathematically sound. It uses 4 features: It calculates single-digit carry-one values $(ST_{n}$), combining them into multidigit carry-one values $(SV_{n}$). Any $SV_{n}$ uncertain (U) values are refined to 0 or 1 over tokens (highlighted). By the “+” token, $SV2$ gives the A3 value as 0 or 1. The other answer digits $A_{n}$ are calculated from $SA_{n}$ and $SV_{n-1}$ values.
We construct (effectively infinite) synthetic training data enriched with the hardest edge cases, and train 49 small transformer models (2–3 layers, 3–4 attention heads) on 5 to 12-digit arithmetic. These models converge rapidly, reaching >>99.999% accuracy. They succeed on the cascading carry one and borrow one cases where LLMs fail e.g. 555555555 + 444444448 = 1000000003. Our largest model has  10M parameters—over four orders of magnitude smaller than GPT-3—and trains to near-perfect accuracy in under an hour on a single GPU, providing an efficient, reproducible, and interpretable testbed for mechanistic analysis.
Figure 2: We score the addition and subtraction capability of 180 public LLMs. A score of 5 means the LLM handled two 5 digits numbers correctly but failed with 6 digit numbers. 7% of addition and 12% of subtraction models get the maximum test score 15. The top models can call external tools.
We introduce a novel, simple, mathematically exact algorithm for n-digit addition (Figure 1). The algorithm respects the transformer’s left-to-right processing constraint, handling high-value digits before low-value ones. We also introduce a novel, exact algorithm for subtraction.
We then ask whether our trained models actually implement these algorithms. To do so, a model must realize each algorithmic subtask. We hypothesize the existence of specific calculation “nodes” capable of carrying out these subtasks, and develop code to automatically search for candidate nodes. We evaluate these candidates through systematic ablations. Each ablation is designed around a specific subtask, allowing us to precisely predict how the model’s output should change when we intervene. The predictions match perfectly: intervening on the identified node produces exactly the expected effect. Every model we study contains nodes corresponding to all subtasks, though in some cases a single logical role is distributed across two attention heads.
The algorithms also impose strict constraints on node ordering and relationships. For example, certain nodes must appear earlier in the computation sequence, while others depend on them downstream (see Figure 1). We test these constraints and find they are consistently satisfied across seeds, model sizes, and digit ranges. Together, these results show that all our accurate models converge on the same algorithmic solution.
To make this analysis reproducible, we release a general-purpose interpretability library. https://github.com/PhilipQuirke/quanta_mech_interp. It provides tools for (1) node characterization, search, and ablation, (2) systematic testing of node relationships and constraints, and (3) visualizations. It is a foundational tool used by this project https://github.com/PhilipQuirke/quanta_maths
Our contributions are fivefold: (i) a large-scale survey of LLM arithmetic capabilities, (ii) enriched datasets for addition and subtraction, (iii) 49 small transformer models trained to near-perfect accuracy, (iv) human-comprehensible exact addition and subtraction algorithms and (v) a reusable interpretability toolkit demonstrating that these models implement the algorithms.
We conclude that transformers can implement exact n-digit addition and subtraction through interpretable circuits. With appropriate training and data, larger LLMs could in principle adopt the same mechanisms.
@@ 2 Related Work
Mechanistic understanding of arithmetic in transformers. Early work by Nanda et al. (2023) revealed that one-layer models implement modular addition through discrete Fourier transforms, converting addition into rotations in frequency space - showing transformers can find unintuitive algorithms for basic mathematics. Quirke and Barez (2024) detailed an algorithm for n-digit addition in a single-layer model that achieved 99% accuracy, identifying a cascading carry-one failure mode. Subsequent work on training dynamics (Musat, 2024) and the finding that even random transformers exhibit algorithmic capabilities (Zhong and Andreas, 2024) suggest architectural biases toward systematic computation.
Architectural modifications and specialized models. Several approaches achieve near-perfect arithmetic accuracy through architectural changes. Reformatting inputs to present least-significant digits first by adding position embeddings and zero-padding strategies (McLeish et al., 2024). Cho et al. (2024) align computational flow with carry propagation. Specialized models demonstrate superior performance: MathGLM’s 10M parameter model achieves 100% accuracy on integer addition through step-by-step reasoning (Yang et al., 2023), while Qiu et al. (2024) reached 99.9% accuracy on 5-digit multiplication with a tiny transformer that outperforms GPT-4. Subtraction is much less studied, though Zhang et al. (2024) identified symmetries between addition and subtraction circuits, with the same attention heads handling both operations.
Production systems versus specialized models. The Claude 3.5 Haiku circuit analysis Lindsey et al. (2025) reveals how a frontier model implements addition. Rather than the clean n-digit algorithms found in specialized models, Claude employs a “bespoke” circuit combining lookup tables for memorized facts with magnitude estimation pathways. This distributed circuit achieves perfect accuracy on 2-digit addition using mechanisms distinct from both human methods and specialized models. The gap between Claude’s complex implementation on short addition questions and the elegant n-digit algorithms in specialized models illustrates a fundamental challenge: production LLMs trained on diverse tasks develop hybrid strategies rather than the interpretable, generalizable algorithms that specialized models achieve.
@@ 3 Methodology
Transformer models learn addition algorithms different from traditional human methods. We define an alternative, mathematically-equivalent framework for addition and demonstrate our models implement this approach.
@@ 3.1 Mathematical Framework
For addition and subtraction of two $n$-digit numbers, we use the Fig. 3 notation. (Detail in App. A).
@@
Figure 3: For 5-digit addition and subtraction, our notation for the main input tokens is $D4$, …, $D0$ and $D^{\prime}4$, …, $D^{\prime}0$. For output tokens it is $A6$, .., $A0$. For n-digits, we use the notation $D_{n}$, .., $D0$, $D^{\prime}_{n}$, .., $D^{\prime}0$ and $A_{n}$, .., $A0$.
3.2 Addition Algorithm Description
For addition, aligned with Quirke and Barez (2024), we define the “Base Add” subtask $SA_{n}$ to compute the digit-wise sum modulo 10:
$SA_{n}=(D_{n}+D^{\prime}_{n})\mod 10$ (1)
@@ 3.2 Addition Algorithm Description
To solve $n$-digit addition with high accuracy, the framework must handle carry bits that cascade through multiple digits. For an autoregressive model to correctly predict the first answer digit in “555+448=+1003”, as “1”, it must cascade the carry bit from the rightmost digit to the leftmost digit in a single forward pass. This is particularly challenging for an autoregressive model that processes tokens from left to right. Aligned to this challenge we define two new subtasks.
First, we introduce subtask ST. It classifies a digit pair sum as definitely causing a carry (e.g. 6+7), definitely not causing a carry (e.g. 2+3), or possibly causing a carry (e.g. 5+4). The 5+4 case is uncertain as the digits sum to 9 and a carry from the next-lower-value digit would cause a carry:
[…]
The “TriAdd” function combines two $ST_{n}$ style values X and Y. Only if both X and Y are U, does TriAdd return U (there is still uncertainty), otherwise it returns 0 or 1 (the carry bit value is known):
[…]
Finally, we introduce the SV subtask that uses TriAdd to handle multidigit carry cascades. We describe $SV_{n}$ using the first 3 examples:
$SV0$ $=ST0$ (4)
$SV1$ $=\text{TriAdd}(ST1,ST0)$ (5)
$SV2$ $=\text{TriAdd}(\text{TriAdd}(ST2,ST1),ST0)$ (6)
The highlighted (solid-color) part of Figure 1 illustrates the SV2 calculation. Note that the calculation first combines ST2 and ST1 (the higher value digits) before combining the result with ST0. This formulation mirrors the left to right order in which the model processes tokens i.e. from higher value digits to lower value digits. The last term in all $SV_{n}$ calculations is ST0 which is always 0 or 1, so all $SV_{n}$ evaluate 0 or 1. That is, at the end of the $SV_{n}$ calculation any uncertainty has been resolved.
The $SV_{n}$ are perfectly accurate multidigit cascade carry values. Their calculation occurs across multiple tokens and is completed by the “=” token. Combining the $SV_{n}$ values with the $SA_{n}$ values as shown in Figure 1 gives perfectly accurate answer tokens.
3.3 Subtraction Algorithm Description
Our subtraction algorithm mirrors our addition algorithm, replacing addition’s “cascading carry one” with subtraction’s “cascading borrow one”. Both operations use the same iterative approach to reducing “cascading” uncertainty token by token.
We introduce “Base Diff” $MD_{n}$ defined as $D_{n}$- $D^{\prime}_{n}$ modulo 10 (paralleling $SA_{n}$). We also introduce $MB_{n}$ for the single-digit “borrow one” case (paralleling $ST_{n}$):
[…]
Finally, we introduce $MV_{n}$ which parallels $SV_{n}$ to handle the “cascading borrow one” edge case. With this formulation, the addition and subtraction algorithms have the same structure (just replacing $SA_{n}$ with $MD_{n}$, $ST_{n}$ with $MB_{N}$, and $SV_{n}$ with $MV_{N}$).
Subtraction poses an additional algorithmic difficulty - some questions give a positive answer (e.g. 325-129=+196) and others a negative answer (e.g. 325-329=-004). Note that all the answer digits are different in the above examples. While there are multiple ways this could be handled mathematicaly, we introduce “Neg Diff” $ND_{n}$ defined as $D^{\prime}_{n}$- $D_{n}$ modulo 10 (the opposite of $MD_{n}$). At the “=” token, the $MV_{n}$ avalues are known, and the algorithm selects $MD_{n}$ or $MD_{n}$ values as appropriate.
Figure 4: Our subtraction algorithm parallels our addition algorithm but it refines “cascading borrow one” uncertainty over multiple tokens. The refined $MV2$ value determines both the answer sign (“+” or “-”) and which digit values ($MD_{n}$ or $ND_{n}$) to use for the final output.
In Fig. 4, the value $MV2$ determines the answer sign. While the following calculations could use $MV2$ too, in fact they attend to the answer sign itself. We introduce the subtask SGN to reflect this.
3.4 Mixed Algorithm Description
For mixed models handling both operations, we add the subtask OPR meaning “attends to the prompt operation token”. The mixed models uses OPR and SGN to determine whether to output addition, positive-subtraction, or negative-subtraction answer digits.
@@ 4 Experiments
4.1 Training Models
Our models’ vocabulary consisted of 14 tokens: the digits 0-9 plus the mathematical operators +, -, =, *, and /. (Also see App. C). We created an infinite training dataset enriched with rare numerical combinations (mainly cascading use cases). (Also see App. D)
We trained addition-only, subtraction-only, and mixed (addition and subtraction) models. We systematically explored different model architectures and settled on 2 layers and 3 or 4 attention heads. (Also see App. F)
We trained used a batch size of 64, learning rate of 0.00008, and weight decay of 0.1. The loss function was defined as the mean negative log likelihood across all output tokens. Training was stopped when the loss was very low e.g. $2\times 10^{-8}$. (Also see App. G, Fig. 5, Fig. 9).
Each model’s accuracy was tested on 1 million questions. Robust statistical accuracy ranges were calculated, but for ease of presentation, we say a model has 99.999% accuracy if it got less than 10 of the 1 million questions wrong. Most models achieved this. The resulting 49 models, including detailed accuracy results, are summarized in Tab. 5 and Tab. 6.
Figure 5: The 5- to 15-digit 2-layer 3-head addition models have very low loss. With more digits training takes longer. Details in Tab.5.
4.1.1 Training mixed models
We trained 14 mixed models with various architectures, including 6- and 10-digit models with 2-3 layers and 3-4 attention heads (see Tab.6). Each model was initialized using weights from a highly accurate addition-only model, specifically copying the weights into the first two layers and first three heads of the mixed model. During training, we used a curriculum of 80% subtraction and 20% addition examples, using enriched data for both operations (See Fig. 6).
This initialization approach proved effective, with most models quickly achieving high accuracy on both operations. We explored alternative training strategies, such as periodically “freezing” the inserted attention heads or MLPs by restoring their original addition-model weights every 100 training steps. However, these more constrained approaches resulted in lower accuracy, suggesting that some flexibility in modifying the initialized weights was necessary for optimal performance (Also see App. K)
Figure 6: Mixed model initialization and analysis: (1) Train an accurate 5-digit addition model. (2) Reverse-engineer it to identify the implemented algorithm. (3) Insert its attention head and MLP weights (in brown) into a larger model. (4) Train the new model on 80% subtraction and 20% addition. (5) The resulting model predicts accurately and reuses the inserted addition circuits for both tasks.
We did a seed sensitivity analysis. App. L includes a visualization Fig. 15 of the range of training losses for different model categories.
4.2 Experimental Results
To verify whether our models implement the algorithms described above, we developed a systematic framework for analyzing model internals. For each trained model, we investigated whether it contains the specific computational components our algorithms require and whether these components interact according to the algorithmic constraints.
4.2.1 Implementation Analysis Framework
Consider a 6-digit addition model. If it implements our algorithm, it must satisfy these constraints:
  • Functional accuracy: The model must correctly add any two 6-digit numbers, selecting the unique correct answer from nearly 2 million possible answers.
  • Complete component coverage: The model must implement the required subtasks for each digit position i.e. 6 $SA$, 6 $ST$, 6 $SV$, and 7 answer digit finalization nodes.
  • Component-level behavior: Each node implementing a subtask must obey certain conditions. For instance, a node implementing $ST2$ must:
    • Attend primarily to tokens $D_{2}$ and $D^{\prime}_{2}$ (post-softmax attention >0.01>0.01)
    • Be positioned after $D^{\prime}_{2}$ but before the “=” token
    • A Primary Component Analysis (PCA) should show 3 distinct clusters. (See Fig. 7.)
    • Given three test question sets, created specifically to trigger the 3 $ST2$ values, each question set should align with exactly one of the PCA clusters.
  • Algorithmic ordering constraints: Node locations must respect computational dependencies. Since SV1 = TriAdd(ST1, ST0), a node computing SV1 must be positioned after nodes computing ST1 and ST0. Our 6-digit model has over 30 such ordering constraints.
  • Causal verification through intervention: When we identify a candidate node for a specific subtask (e.g. ST2), we can construct paired test questions differing only in that subtask’s value. By intervening on the candidate node and swapping its activations between the test pairs, we can predict the model’s output from nearly 2 million possible answers.
For models with more than 6 digits, the constraints list grows further. The subtraction and mixed algorithms each have a more extensive constraints list. With 49 models to test, we developed an automated testing framework to:
  • Systematically search for nodes over each subtask and digit position combination retaining only those that satisfy all the constraints in our algorithm
  • Compare the list of nodes found with the algorithmic requirements (including intra-node relationships).
  • Visualize the results.
4.2.2 Addition model results
We tested a single 1-layer addition model. Our framework found the $SA_{n}$ and $SC_{n}$ (single-digit carry calculated during answer tokens) nodes as described by Quirke and Barez (2024) that give the model its 99% accuracy.
Testing our fifteen 2-layer addition models, we find all models contain the subtasks required by our addition algorithm, and satisfy the algorithmic requirements. This consistent pattern suggests we have identified a robust and generalizable algorithm that naturally arises when transformer models are trained on addition in our setup. There is some variability between models:
  • In a given model, $ST_{N}$ is not as “tidy” as shown in Fig. 1. Different models select different attention heads to calculate say ST2. Redundantly, some models have two nodes that both calculate say ST1. See Tab. 1.
  • Some models combine 2 attention heads (with the same token position and layer) to calculate say SA2. Many models only use one attention head.
  • For ST nodes, the PCA trigrams have difference appearances in each attention head (- but all have 3 clusters aligned to the ST values. See Fig. 7.
  • All models have $ST$ nodes, but some models also have and use $SC$ nodes. The $SC$ nodes are redundant: the $ST$ values are equivalent and the $SV$ values are superior! See Tab. 1.
Figure 7: For a 5-digit addition model, these three sample PCAs each show 1 attention head at a token position tested for 1 answer digit. Each dot represents a question, colored by the question’s expected ST value (0, 1, or U). While the PCA data varies across cases, the 3 ST classes form visually distinct clusters in each plot, suggesting that the attention head’s activation patterns differ systematically based on ST value and that the model could potentially distinguish between these 3 cases.
Table 1: For a sample 5-digit addition model, we show a compacted location map for the SASA, SCSC and STST subtasks. Interestingly 1) the $ST$ nodes are in a semi-random order 2) the second $ST1$ node is redundant 3) the model uses SCSC nodes which are redundant 4) each $SA$ subtask is shared across two attention heads.